OpenStudy (anonymous):

a measure of the quantity of bacteria in a cubic centimeter of blood of an infected person is modeled by Q(t) = 80t(.92^t) t is the time in days since the person became ill. how fast is the quantity changing 7 days after the onset of the illness?

OpenStudy (jack1):

if Q(t) = 80t(.92^t) and t = time and time = 7days sub t=7 into Q(t) = 80t(.92^t) becomes ====> Q(7) = 80 x 7 x (.92^7)\[Q(7) = 80 \times 7 \times (.92^7)\]

OpenStudy (jack1):

can u solve from here? @wittless29 ?

OpenStudy (anonymous):

that's not what the answer is on the key, I wish it was that straightforward. thanks for trying though

OpenStudy (jack1):

so is the equation i posted above matching ur's in the book/ pdf?

OpenStudy (jack1):

or is it mor like: \[Q(7) = 80 \times 7 ^{ (0.92^7)}\]?

OpenStudy (jack1):

or something else?